OpenStudy (ranya99):

A crate with mass 32.5 kg initially at rest on a warehouse floor is acted on by a net horizontal force of 140. N. What acceleration is produced? How far does the crate travel in 10.0 s?

OpenStudy (naveenbabbar):

a= Net force/mass s=(1/2) x acceleration x square of time

rvc (rvc):

Newton's Second law F=ma so a= f/m now since initial velocity u=0 n we want distance i.e s s=ut+1/2 at^2 here u=0 a= the value u obtain m/s^2 t= 10 sec thats it :)

rvc (rvc):

@Ranya99 any doubts?

OpenStudy (ranya99):

@rvc no, no doubts. Got it, thanks! :)

rvc (rvc):

okay enjoy learning All d best :)

OpenStudy (ranya99):

@rvc Can you help me with another problem?